A uniform steel plate has an area of 0.819 m2 When subjected to a temperature difference between its sides, a heat current of 25.7 kW is found to flow temperature difference when the plate is 5.73 cm thick? The thermal conductivity of steel is 50.2 WI(m K). it. What is the temperature gradient? What is the Temperature gradient: Number K/ m Temperature difference: Number () The heat current is the amount of heat that flows per unit time
Answer #1

Write the expression of heat flow through a conductor -

Q = k*A*ΔT / t --------------------------------------------------(i)

Where -

Q = 25.7 kW = 25.7 x 1000 = 25700 W

A = 0.819 m^2

ΔT = ?

t = 5.73 cm = 0.0573 m

k = 50.2 W/m·K

Put these values in expression (i) -

25700 W = 50.2W/m·K * 0.819m² * ΔT / 0.0573m

=> ΔT = (25700 * 0.0573) / (50.2 * 0.819) =  35.8ºC

So, temperature difference between the sides of the plate = 35.8ºC (Answer of second part)

Temperature gradient = ΔT/t = 35.8 / 0.0573 = 624.8 K/m (Answer of first part)